# multimotion estimation problem MOT multiobject tracking

### v3

@@ 3.2 Graph Construction
where $\bar{d}_{ij}$ is the mean distance between $\mathscr{p}^{i}$ and $\mathscr{p}^{j}$. The variance between tracklets moving similarly is small, making this edge cost small for tracklets on the same rigid body. The graph consists of vertices for all tracklets and the $k$ least-costly edges from each vertex. This $k$-nearest neighbors graph is unlikely to connect independently moving pairs of tracklets so it reasonably and efficiently approximates the rigid-body assumption. Its connectivity forms the basis for label generation (Section 3.3) and assignment (Section 3.4), and an over- or under-connected graph can lead to an under- or overfit solution, respectively.
@@ 3.3 Label Proposal
The motions in the scene are segmented by assigning tracklets to a finite set of motion labels, $\mathscr{L}$. This set must be updated regularly because dynamic environments are constantly changing. New labels are proposed by splitting existing labels whenever their motions could be explained by multiple trajectories. A potential new label, $\ell{}^{\prime}$, is generated for each fully-disjoint component of the subgraph, $\mathcal{N}_{\ell{}}\subseteq\mathcal{N}{}$, defined by the current label support, $\mathscr{P}_{\ell{}}$. This prioritizes graph smoothness while allowing large label subgraphs consisting of distinct rigid motions to be split into new labels.
@@ 3.3.1 Estimating New Labels
Each label is estimated as if it corresponds to the static portion of the scene using standard egomotion techniques. The new egomotion hypothesis, ${{}^{\ell{}^{\prime}}\mathscr{T}_{C}}$, is calculated from the dominant motion of the corresponding subgraph component using frame-to-frame RANSAC, which segments the inliers and outliers of that motion (Fischler and Bolles, 1981). The frame-to-frame RANSAC procedure estimates the observed motion between pairs of consecutive frames, ${{}^{\ell{}^{\prime}}\mathbf{T}_{{{C_{k}}{C_{k-1}}}}}$, by sampling three tracklets in both frames and assuming they are static. The transform is estimated in a blockwise manner according to (1). The translational component is calculated by
@@
less than a chose threshold, $e_{\mathrm{th}}$, where $\mathbf{s}\left(\cdot\right)$ applies the sensor projection (e.g., nonlinear stereo camera; Appendix B). The process of sampling, estimation, and evaluation is iterated a chosen number of times, $N_{\rm{RSAC}}$, and the transform with the most inliers is kept. This process is repeated independently for each pair of consecutive frames in the estimation window to generate a new trajectory for the label. This frame-to-frame procedure is used to propose a new trajectory for each extant label. Any tracklets found to be outliers of the newly estimated trajectories are appended to the outlier label, $\mathscr{O}{}$. After the outliers are removed from all existing label sets, new labels are generated from the outlier label using the same proposal process.
@@ 3.4 Label Assignment
where $\ell{}\left(\mathscr{p}\right)$ gives the label currently assigned to $\mathscr{p}$ and $\lambda_{\mathrm{sm}}$ is a user-selected proportionality parameter. Increasing the proportionality parameter can lead to a spatially smoother solution at the cost of ignoring small, but clearly distinct, motions. Decreasing the parameter can isolate small motions but can also lead to more spurious motion detections. The residual and smoothness portions of the energy are minimized using the convex relaxation algorithm (CORAL; Amayo et al., 2018), a multimodel-fitting algorithm that relaxes discrete labeling to allow for efficient parallelization on GPGPU devices. This “soft” labeling assigns each tracklet a score for how applicable each label is to the observed data. The score is in the continuous range $\left[0,1\right]$ and the scores for each tracklet across all labels sum to $1$. The soft labels are discretized by taking the strongest label for each point. Points with ambiguous soft labels (e.g., less than $0.5$) are labeled as outliers to avoid mislabeling points. The total energy is finally minimized by merging similar labels to avoid oversegmentation and overfitting (Section 3.4.4).
@@ 3.4.1 Residual Cost
where $e_{k}$ is defined in (4) and $t_{1}$ and $t_{2}$ are the first and last time stamps that the tracklet is observed, respectively. Maximum error penalizes tracklets that do not fit a motion, even for a single frame. This conservative labeling helps preserve the accuracy of the trajectory estimation. The residual cost of the outlier label, $\mathscr{O}{}$, is uniquely defined to have low cost for points not well explained by existing labels. The cost decays exponentially with the best-fitting label cost,
@@ 3.4.4 Label Merging
Label merging reduces the total energy in (5) by addressing the model complexity term. Two labels, $\ell{}$ and $\ell{}^{\prime}$, are merged if relabeling all tracklets from one label, $\mathscr{P}_{\ell{}^{\prime}}$, to the other label, $\ell{}$, would decrease the total energy. This occurs when the increase in residual error is less than the cost of using the label and any change in smoothness. When more than one merge would reduce the total energy, the merge that results in the greatest decrease in cost is chosen. Merging continues until no more merges would reduce the total energy in (5). The outlier label, $\mathscr{O}$, is excluded from merging. The label splitting, assignment, and merging stages (Sections 3.3 and 3.4) are iterated until the labels converge or a maximum number of iterations, $N_{\mathrm{conv}}{}$, have been reached. Label convergence can be defined relative to various criteria, such as the total energy, the change in energy, or the change in the labeling.
@@ 4 Batch $SE\left(3\right)$ Estimation
The $SE\left(3\right)$ trajectories of the sensor and each object in the scene are estimated from the segmented tracklets. Batch trajectory estimation techniques traditionally calculate the motion of a sensor relative to a set of points assuming they are static. This is appropriate for egomotion estimation (Section 4.1), but third-party trajectory estimation requires extending these techniques to include the geocentric (i.e., quasi-inertial) frame defined by the egomotion estimation (Section 4.2). Initially, each segmented motion represents an egomotion trajectory hypothesis, ${{}^{\ell{}}\mathbf{T}_{{{C_{k}}{C_{k-1}}}}}$, under the assumption that the associated tracklets, $\mathscr{P}_{\ell{}}$, are static. Once the true static label is identified, its corresponding egomotion estimate can be used to estimate the remaining third-party motions in a geocentric frame. These geocentric trajectories can be estimated without making any assumptions about the object or sensor motion other than rigid motion. The estimation can be made more robust by using more expressive motion models that constrain the trajectories to the real-world kinematics of the scene. Examples of generalized, physically founded $SE\left(3\right)$ motion models include the WNOA prior described by Anderson and Barfoot (2015) and the WNOJ prior described by Tang et al. (2019). The WNOA prior extends the pose-only trajectory state to also include velocity and the WNOJ prior extends it to include both velocity and acceleration. The applicability of any motion model depends on how well it matches the motions in an environment. The pose-only estimator (Sections 4.1.1 and 4.2.1) can estimate complex rigid motions but can also generate physically implausible motions. The pose-velocity (Sections 4.1.2 and 4.2.2) and pose-velocity-acceleration (Sections 4.1.3 and 4.2.3) estimators penalize deviation from a constant body-centric velocity or acceleration, respectively. They are generally applicable because objects tend to move smoothly through their environment, but they are not robust to motions that deviate significantly from these assumptions. Each estimator brings advantages and disadvantages in the context of the MEP, and this investigation shows how MVO can leverage different estimators in different applications (Sections 6 and 7).
@@ 4.1 Egomotion Estimation
The egomotion label, $C$, may be selected by a variety of heuristics. It can be chosen as the label with the largest support, as in VO, or with the most similar motion to the previous egomotion estimate. This can be combined with external sensors or application-specific heuristics, such as attention masks that prioritize parts of the field of view that usually contain static background objects, to robustly identify the true egomotion label. Once identified, the point measurements in the egomotion label are assumed to be static. The trajectory is therefore estimated relative to a geocentric frame and inertial motion priors can be used in the estimation.
@@ 4.1.1 Pose-Only Estimator
A discrete single-motion bundle adjustment (e.g., Barfoot, 2017) defines the egomotion state, $\mathbf{x}{}$, to include both the estimated pose transforms, $\mathscr{T}_{C}\coloneqq\left({\mathbf{T}_{C_{k}C_{1}}}\right)_{k=2\dots K}$, and the landmark points, $\left\{\mathbf{p}^{{j_{1}}{C_{1}}}_{C_{1}}\right\}_{j=1\dots|\mathscr{P}_{C}|}$. The state $\mathbf{x}{}_{jk}\coloneqq\left\{{\mathbf{T}_{C_{k}C_{1}}},\mathbf{p}^{{j_{1}}{C_{1}}}_{C_{1}}\right\}$ is defined for each pair of transforms and points in label $C$, i.e., $\mathbf{T}_{C_{k}C_{1}}\in\mathscr{T}_{C}$ and $\mathbf{p}^{{j_{1}}{C_{1}}}_{C_{1}}\in\mathscr{p}^{C}$. The batch size, $K$, is the length of the estimation window, which can either be the entire sequence, or some fixed-length sliding window (Section 5). Each observation, $\mathbf{y}_{jk}$, of point, $\underrightarrow{p}^{j}$, at pose, ${\mathbf{T}_{C_{k}C_{1}}}$, is modeled as
@@
is the residual between the observed and modeled tracklet position. This cost is linearized about an operating point, $\mathbf{x}{}_{\mathrm{op}}$, and then minimized using Gauss-Newton. The linearized cost function is
@@
$\frac{\partial\mathbf{z}}{\partial\delta\mathbf{x}{}}\bigg{\rvert}_{\mathbf{x}{}_{\mathrm{op},jk}}$ $=\begin{bmatrix}\left({\mathbf{T}_{\mathrm{op},{C_{k}}{C_{1}}}}\boldsymbol{p}^{j_{1}C_{1}}_{\mathrm{op},C_{1}}\right)^{\odot}&{\mathbf{T}_{\mathrm{op},{C_{k}}{C_{1}}}}\mathbf{D}\end{bmatrix},$ (10)
$\mathbf{D}$ $=\begin{bmatrix}\mathbf{1}{}\\ \mathbf{0}{}^{T}\end{bmatrix},$ (13)
@@
The state perturbation, $\delta\mathbf{x}{}$, consists of the transform perturbations, $\{\delta\boldsymbol{\xi}_{k}\in\mathbb{R}^{6}\}$, and landmark perturbations, $\{\delta\boldsymbol{\zeta}_{j}\in\mathbb{R}^{3}\}$. The indicator matrix $\mathbf{P}_{jk}$ is defined such that $\delta\mathbf{x}{}_{jk}=\mathbf{P}_{jk}\delta\mathbf{x}{}$ is the perturbation of the $\left\{\mathbf{T}_{\mathrm{op},C_{k}C_{1}},\mathbf{p}^{j_{1}C_{1}}_{\mathrm{op},C_{1}}\right\}$ pair. The optimal perturbation, $*\delta\mathbf{x}{}^{*}$, for minimizing the linearized cost function is the solution to the linear equation $\mathbf{A}\delta\mathbf{x}{}^{*}=\mathbf{b}$. Each element of the state is then updated according to
@@ 4.1.2 Pose-Velocity Estimator
The pose-only estimator is applicable to any rigid-body motion, but is not constrained to physically plausible motions. This can be mitigated by constraining the trajectory to a locally constant velocity. The pose-velocity estimator employs an $SE\left(3\right)$ WNOA prior (Anderson and Barfoot, 2015),
@@
which defines the trajectory acceleration, $\dot{\boldsymbol{\varpi}}_{C}$, as a continuous zero-mean, white noise Gaussian process with power spectral density matrix, $\mathbf{Q}_{c}\in\mathbb{R}^{6\times 6}$. The prior models the velocity as being locally constant, but can vary if there are sufficient observations to support it. The pose-velocity estimator state, $\mathbf{x}{}$, comprises the estimated pose transforms, $\mathscr{T}_{C}$, body-centric velocities, $\left(\boldsymbol{\varpi}_{C_{k}}\right)_{k=1\dots K}$, and the labeled landmark points, $\left\{\mathbf{p}^{{j_{1}}{C_{1}}}_{C_{1}}\right\}_{j=1\dots|\mathscr{P}_{C}|}$. The state $\mathbf{x}{}_{k}\coloneqq\left\{\mathbf{T}_{C_{k}C_{1}},\boldsymbol{\varpi}_{C_{k}}\right\}$ is defined as the transform and velocity at each time, $t_{k}$, and $\mathbf{x}{}_{jk}\coloneqq\left\{\mathbf{x}{}_{k},\mathbf{p}^{{j_{1}}{C_{1}}}_{C_{1}}\right\}$ is defined for each pair of time steps and points in the label. The prior is characterized by the transition function,
@@
where $\Delta t_{\tau,k}=\tau-t_{k}$. The transition function propagates the local prior state, ${\boldsymbol{\gamma}}_{C_{k}}$, from time $t_{k}$ to time $\tau$ according to
@@
$\boldsymbol{\gamma}_{C_{k}}\left(\tau\right)$ $\coloneqq\begin{bmatrix}\mathbf{T}_{C_{k}}\left(\tau\right)\\ \boldsymbol{\varpi}_{C_{k}}\left(\tau\right)\end{bmatrix}.$ (20)
@@
and $\mathbf{G}_{jk}$ and $\mathbf{E}_{k}$ are the Jacobians of the measurement and prior error functions, respectively. The Jacobian of the measurement model, $\mathbf{G}_{jk}$, is defined according to (8) with the updated motion model Jacobian,
@@
using the $SO\left(3\right)$ lift operator, $\left(\cdot\right)^{\wedge}\colon\mathbb{R}^{3\times 1}\to\mathbb{R}^{3\times 3}$, which is equivalent to the skew-symmetric operator, $\left(\cdot\right)^{\times}$. The adjoint of $\mathbf{T}_{C_{k+1}C_{k}}$ is defined as
@@
The operating point is perturbed according to the transform perturbations, $\{\delta\boldsymbol{\xi}_{k}\in\mathbb{R}^{6}\}$, velocity perturbations $\{\delta\boldsymbol{\varpi}_{k}\in\mathbb{R}^{6}\}$, and landmark perturbations, $\{\delta\boldsymbol{\zeta}_{j}\in\mathbb{R}^{3}\}$, which are stacked to form the full state perturbation, $\delta\mathbf{x}{}$. The indicator matrices $\mathbf{P}_{jk}$ and $\mathbf{P}_{k}$ are defined such that $\delta\mathbf{x}{}_{jk}=\mathbf{P}_{jk}\delta\mathbf{x}{}$ is the perturbation of the $\left\{\mathbf{T}_{\mathrm{op},C_{k}C_{1}},\boldsymbol{\varpi}_{\mathrm{op},C_{k}},\mathbf{p}^{j_{1}C_{1}}_{\mathrm{op},C_{1}}\right\}$ tuple and $\delta\mathbf{x}{}_{k}=\mathbf{P}_{k}\delta\mathbf{x}{}$ is the perturbation of the $\left\{\mathbf{T}_{\mathrm{op},C_{k}C_{1}},\boldsymbol{\varpi}_{\mathrm{op},C_{k}}\right\}$ pair. The optimal perturbation, $*\delta\mathbf{x}{}^{*}$, to minimize the linearized cost is the solution to the linear equation, $\mathbf{A}\delta\mathbf{x}{}^{*}=\mathbf{b}$. Each element of the operating point is then updated using (14) and

### v4

@@ 3.2 Graph Construction
where $\bar{d}_{ij}$ is the mean distance between $\mathscr{p}^{i}$ and $\mathscr{p}^{j}$. The variance between tracklets moving similarly is small, making this edge cost small for tracklets on the same rigid body.
The graph consists of vertices for all tracklets and the $k$ least-costly edges from each vertex. This $k$-nearest neighbors graph is unlikely to connect independently moving pairs of tracklets so it reasonably and efficiently approximates the rigid-body assumption. Its connectivity forms the basis for label generation (Section 3.3) and assignment (Section 3.4), and an over- or under-connected graph can lead to an under- or overfit solution, respectively.
@@ 3.3 Label Proposal
The motions in the scene are segmented by assigning tracklets to a finite set of motion labels, $\mathscr{L}$. This set must be updated regularly because dynamic environments are constantly changing. New labels are proposed by splitting existing labels whenever their motions could be explained by multiple trajectories.
A potential new label, $\ell{}^{\prime}$, is generated for each fully-disjoint component of the subgraph, $\mathcal{N}_{\ell{}}\subseteq\mathcal{N}{}$, defined by the current label support, $\mathscr{P}_{\ell{}}$. This prioritizes graph smoothness while allowing large label subgraphs consisting of distinct rigid motions to be split into new labels.
@@ 3.3.1 Estimating New Labels
Each label is estimated as if it corresponds to the static portion of the scene using standard egomotion techniques. The new egomotion hypothesis, ${{}^{\ell{}^{\prime}}\mathscr{T}_{C}}$, is calculated from the dominant motion of the corresponding subgraph component using frame-to-frame RANSAC, which segments the inliers and outliers of that motion (Fischler and Bolles, 1981).
The frame-to-frame RANSAC procedure estimates the observed motion between pairs of consecutive frames, ${{}^{\ell{}^{\prime}}\mathbf{T}_{{{C_{k}}{C_{k-1}}}}}$, by sampling three tracklets in both frames and assuming they are static. The transform is estimated in a blockwise manner according to (1).
The translational component is calculated by
@@
less than a chose threshold, $e_{\mathrm{th}}$, where $\mathbf{s}\left(\cdot\right)$ applies the sensor projection (e.g., nonlinear stereo camera; Appendix B).
The process of sampling, estimation, and evaluation is iterated a chosen number of times, $N_{\rm{RSAC}}$, and the transform with the most inliers is kept. This process is repeated independently for each pair of consecutive frames in the estimation window to generate a new trajectory for the label.
This frame-to-frame procedure is used to propose a new trajectory for each extant label. Any tracklets found to be outliers of the newly estimated trajectories are appended to the outlier label, $\mathscr{O}{}$. After the outliers are removed from all existing label sets, new labels are generated from the outlier label using the same proposal process.
@@ 3.4 Label Assignment
where $\ell{}\left(\mathscr{p}\right)$ gives the label currently assigned to $\mathscr{p}$ and $\lambda_{\mathrm{sm}}$ is a user-selected proportionality parameter. Increasing the proportionality parameter can lead to a spatially smoother solution at the cost of ignoring small, but clearly distinct, motions. Decreasing the parameter can isolate small motions but can also lead to more spurious motion detections.
The residual and smoothness portions of the energy are minimized using the convex relaxation algorithm (CORAL; Amayo et al., 2018), a multimodel-fitting algorithm that relaxes discrete labeling to allow for efficient parallelization on GPGPU devices. This “soft” labeling assigns each tracklet a score for how applicable each label is to the observed data. The score is in the continuous range $\left[0,1\right]$ and the scores for each tracklet across all labels sum to $1$. The soft labels are discretized by taking the strongest label for each point. Points with ambiguous soft labels (e.g., less than $0.5$) are labeled as outliers to avoid mislabeling points. The total energy is finally minimized by merging similar labels to avoid oversegmentation and overfitting (Section 3.4.4).
@@ 3.4.1 Residual Cost
where $e_{k}$ is defined in (4) and $t_{1}$ and $t_{2}$ are the first and last time stamps that the tracklet is observed, respectively. Maximum error penalizes tracklets that do not fit a motion, even for a single frame. This conservative labeling helps preserve the accuracy of the trajectory estimation.
The residual cost of the outlier label, $\mathscr{O}{}$, is uniquely defined to have low cost for points not well explained by existing labels. The cost decays exponentially with the best-fitting label cost,
@@ 3.4.4 Label Merging
Label merging reduces the total energy in (5) by addressing the model complexity term. Two labels, $\ell{}$ and $\ell{}^{\prime}$, are merged if relabeling all tracklets from one label, $\mathscr{P}_{\ell{}^{\prime}}$, to the other label, $\ell{}$, would decrease the total energy. This occurs when the increase in residual error is less than the cost of using the label and any change in smoothness. When more than one merge would reduce the total energy, the merge that results in the greatest decrease in cost is chosen. Merging continues until no more merges would reduce the total energy in (5). The outlier label, $\mathscr{O}$, is excluded from merging.
The label splitting, assignment, and merging stages (Sections 3.3 and 3.4) are iterated until the labels converge or a maximum number of iterations, $N_{\mathrm{conv}}{}$, have been reached. Label convergence can be defined relative to various criteria, such as the total energy, the change in energy, or the change in the labeling.
@@ 4 Batch $SE\left(3\right)$ Estimation
The $SE\left(3\right)$ trajectories of the sensor and each object in the scene are estimated from the segmented tracklets. Batch trajectory estimation techniques traditionally calculate the motion of a sensor relative to a set of points assuming they are static. This is appropriate for egomotion estimation (Section 4.1), but third-party trajectory estimation requires extending these techniques to include the geocentric (i.e., quasi-inertial) frame defined by the egomotion estimation (Section 4.2).
Initially, each segmented motion represents an egomotion trajectory hypothesis, ${{}^{\ell{}}\mathbf{T}_{{{C_{k}}{C_{k-1}}}}}$, under the assumption that the associated tracklets, $\mathscr{P}_{\ell{}}$, are static. Once the true static label is identified, its corresponding egomotion estimate can be used to estimate the remaining third-party motions in a geocentric frame.
These geocentric trajectories can be estimated without making any assumptions about the object or sensor motion other than rigid motion. The estimation can be made more robust by using more expressive motion models that constrain the trajectories to the real-world kinematics of the scene. Examples of generalized, physically founded $SE\left(3\right)$ motion models include the WNOA prior described by Anderson and Barfoot (2015) and the WNOJ prior described by Tang et al. (2019). The WNOA prior extends the pose-only trajectory state to also include velocity and the WNOJ prior extends it to include both velocity and acceleration.
The applicability of any motion model depends on how well it matches the motions in an environment. The pose-only estimator (Sections 4.1.1 and 4.2.1) can estimate complex rigid motions but can also generate physically implausible motions. The pose-velocity (Sections 4.1.2 and 4.2.2) and pose-velocity-acceleration (Sections 4.1.3 and 4.2.3) estimators penalize deviation from a constant body-centric velocity or acceleration, respectively. They are generally applicable because objects tend to move smoothly through their environment, but they are not robust to motions that deviate significantly from these assumptions. Each estimator brings advantages and disadvantages in the context of the MEP, and this investigation shows how MVO can leverage different estimators in different applications (Sections 6 and 7).
@@ 4.1 Egomotion Estimation
The egomotion label, $C$, may be selected by a variety of heuristics. It can be chosen as the label with the largest support, as in VO, or with the most similar motion to the previous egomotion estimate. This can be combined with external sensors or application-specific heuristics, such as attention masks that prioritize parts of the field of view that usually contain static background objects, to robustly identify the true egomotion label.
Once identified, the point measurements in the egomotion label are assumed to be static. The trajectory is therefore estimated relative to a geocentric frame and inertial motion priors can be used in the estimation.
@@ 4.1.1 Pose-Only Estimator
A discrete single-motion bundle adjustment (e.g., Barfoot, 2017) defines the egomotion state, $\mathbf{x}{}$, to include both the estimated pose transforms, $\mathscr{T}_{C}\coloneqq\left({\mathbf{T}_{C_{k}C_{1}}}\right)_{k=2\dots K}$, and the landmark points, $\left\{\mathbf{p}^{{j_{1}}{C_{1}}}_{C_{1}}\right\}_{j=1\dots|\mathscr{P}_{C}|}$. The state $\mathbf{x}{}_{jk}\coloneqq\left\{{\mathbf{T}_{C_{k}C_{1}}},\mathbf{p}^{{j_{1}}{C_{1}}}_{C_{1}}\right\}$ is defined for each pair of transforms and points in label $C$, i.e., $\mathbf{T}_{C_{k}C_{1}}\in\mathscr{T}_{C}$ and $\mathbf{p}^{{j_{1}}{C_{1}}}_{C_{1}}\in\mathscr{p}^{C}$. The batch size, $K$, is the length of the estimation window, which can either be the entire sequence, or some fixed-length sliding window (Section 5).
Each observation, $\mathbf{y}_{jk}$, of point, $\underrightarrow{p}^{j}$, at pose, ${\mathbf{T}_{C_{k}C_{1}}}$, is modeled as
@@
is the residual between the observed and modeled tracklet position.
This cost is linearized about an operating point, $\mathbf{x}{}_{\mathrm{op}}$, and then minimized using Gauss-Newton. The linearized cost function is
@@
$\frac{\partial\mathbf{z}}{\partial\delta\mathbf{x}{}}\bigg{\rvert}_{\mathbf{x}{}_{\mathrm{op},jk}}$ $=\begin{bmatrix}\left({\mathbf{T}_{\mathrm{op},{C_{k}}{C_{1}}}}\boldsymbol{p}^{j_{1}C_{1}}_{\mathrm{op},C_{1}}\right)^{\odot}&{\mathbf{T}_{\mathrm{op},{C_{k}}{C_{1}}}}\mathbf{D}\end{bmatrix},$
$\mathbf{D}$ $=\begin{bmatrix}\mathbf{1}{}\\ \mathbf{0}{}^{T}\end{bmatrix},$
@@
The state perturbation, $\delta\mathbf{x}{}$, consists of the transform perturbations, $\{\delta\boldsymbol{\xi}_{k}\in\mathbb{R}^{6}\}$, and landmark perturbations, $\{\delta\boldsymbol{\zeta}_{j}\in\mathbb{R}^{3}\}$. The indicator matrix $\mathbf{P}_{jk}$ is defined such that $\delta\mathbf{x}{}_{jk}=\mathbf{P}_{jk}\delta\mathbf{x}{}$ is the perturbation of the $\left\{\mathbf{T}_{\mathrm{op},C_{k}C_{1}},\mathbf{p}^{j_{1}C_{1}}_{\mathrm{op},C_{1}}\right\}$ pair.
The optimal perturbation, $\delta\mathbf{x}{}^{*}$, for minimizing the linearized cost function is the solution to the linear equation $\mathbf{A}\delta\mathbf{x}{}^{*}=\mathbf{b}$. Each element of the state is then updated according to
@@ 4.1.2 Pose-Velocity Estimator
The pose-only estimator is applicable to any rigid-body motion, but is not constrained to physically plausible motions. This can be mitigated by constraining the trajectory to a locally constant velocity.
The pose-velocity estimator employs an $SE\left(3\right)$ WNOA prior (Anderson and Barfoot, 2015),
@@
which defines the trajectory acceleration, $\dot{\boldsymbol{\varpi}}_{C}$, as a continuous zero-mean, white noise Gaussian process with power spectral density matrix, $\mathbf{Q}_{c}\in\mathbb{R}^{6\times 6}$. The prior models the velocity as being locally constant, but can vary if there are sufficient observations to support it.
The pose-velocity estimator state, $\mathbf{x}{}$, comprises the estimated pose transforms, $\mathscr{T}_{C}$, body-centric velocities, $\left(\boldsymbol{\varpi}_{C_{k}}\right)_{k=1\dots K}$, and the labeled landmark points, $\left\{\mathbf{p}^{{j_{1}}{C_{1}}}_{C_{1}}\right\}_{j=1\dots|\mathscr{P}_{C}|}$. The state $\mathbf{x}{}_{k}\coloneqq\left\{\mathbf{T}_{C_{k}C_{1}},\boldsymbol{\varpi}_{C_{k}}\right\}$ is defined as the transform and velocity at each time, $t_{k}$, and $\mathbf{x}{}_{jk}\coloneqq\left\{\mathbf{x}{}_{k},\mathbf{p}^{{j_{1}}{C_{1}}}_{C_{1}}\right\}$ is defined for each pair of time steps and points in the label.
The prior is characterized by the transition function,
@@
where $\Delta t_{\tau,k}=\tau-t_{k}$.
The transition function propagates the local prior state, ${\boldsymbol{\gamma}}_{C_{k}}$, from time $t_{k}$ to time $\tau$ according to
@@
$\boldsymbol{\gamma}_{C_{k}}\left(\tau\right)$ $\coloneqq\begin{bmatrix}\mathbf{T}_{C_{k}}\left(\tau\right)\\ \boldsymbol{\varpi}_{C_{k}}\left(\tau\right)\end{bmatrix}.$
@@
and $\mathbf{G}_{jk}$ and $\mathbf{E}_{k}$ are the Jacobians of the measurement and prior error functions, respectively.
The Jacobian of the measurement model, $\mathbf{G}_{jk}$, is defined according to (8) with the updated motion model Jacobian,
@@
using the $SO\left(3\right)$ lift operator, $\left(\cdot\right)^{\wedge}\colon\mathbb{R}^{3\times 1}\to\mathbb{R}^{3\times 3}$, which is equivalent to the skew-symmetric operator, $\left(\cdot\right)^{\times}$.
The adjoint of $\mathbf{T}_{C_{k+1}C_{k}}$ is defined as
@@
The operating point is perturbed according to the transform perturbations, $\{\delta\boldsymbol{\xi}_{k}\in\mathbb{R}^{6}\}$, velocity perturbations $\{\delta\boldsymbol{\varpi}_{k}\in\mathbb{R}^{6}\}$, and landmark perturbations, $\{\delta\boldsymbol{\zeta}_{j}\in\mathbb{R}^{3}\}$, which are stacked to form the full state perturbation, $\delta\mathbf{x}{}$. The indicator matrices $\mathbf{P}_{jk}$ and $\mathbf{P}_{k}$ are defined such that $\delta\mathbf{x}{}_{jk}=\mathbf{P}_{jk}\delta\mathbf{x}{}$ is the perturbation of the $\left\{\mathbf{T}_{\mathrm{op},C_{k}C_{1}},\boldsymbol{\varpi}_{\mathrm{op},C_{k}},\mathbf{p}^{j_{1}C_{1}}_{\mathrm{op},C_{1}}\right\}$ tuple and $\delta\mathbf{x}{}_{k}=\mathbf{P}_{k}\delta\mathbf{x}{}$ is the perturbation of the $\left\{\mathbf{T}_{\mathrm{op},C_{k}C_{1}},\boldsymbol{\varpi}_{\mathrm{op},C_{k}}\right\}$ pair.
The optimal perturbation, $\delta\mathbf{x}{}^{*}$, to minimize the linearized cost is the solution to the linear equation, $\mathbf{A}\delta\mathbf{x}{}^{*}=\mathbf{b}$. Each element of the operating point is then updated using (9) and